OpenStudy (anonymous):

A 10 kg weight is raised at a hight 2m with an acceleration 1.5m/s2, find total work done?

Parth (parthkohli):

http://imgur.com/uV6FCja

OpenStudy (imtiaz7):

By using formula to find work: W=F.S but Acc. to 2nd law of motion: F=ma by putting the value of F we get: W=ma*S By putting the values of mass, acceleration and distance: W=10*1.5*2 W=30J